OpenStudy (anonymous):

Multiply (a+4)^3 I have a^3+12a^2+?a+64 I am having trouble finding the number with the ?a

OpenStudy (anonymous):

ok the answer that i have that is really not completed is correct or not?

OpenStudy (anonymous):

ok i will try i am kinda rusty on this i havent did this for a long time

sam (.sam.):

For (a+4)^3 (a+4)^3=(a+4)(a+4)(a+4) So, you can apply FOIL to (a+4)(a+4)(a+4) , two times. It is not complete, you have to apply the FOIL method, (a+4)(a+4)(a+4) When you distribute, you will get (a^2+8a+16)(a+4) Then try expanding it again.

sam (.sam.):

There are some examples learning FOIL here: http://www.mathsisfun.com/algebra/polynomials-multiplying.html

OpenStudy (anonymous):

a*a*a=a^3 4+4+4=12a

OpenStudy (anonymous):

no.. it is 48a

OpenStudy (anonymous):

48a where i have my ? at

OpenStudy (anonymous):

i am trying to figure out how o get that number thats all and i keep getting answers that is not helping on how to find...

OpenStudy (anonymous):

so could you please show me how did you get 48a?

OpenStudy (anonymous):

\[(a+b)^3=a^3+3a^2b+3ab^2+b^3\]

OpenStudy (anonymous):

Or If u knw Binomial Theorem Use it..)

OpenStudy (anonymous):

ok lets see here i am going to figure out how you got 48a

sam (.sam.):

(a^2+8a+16)(a+4) When you're here, there's a shortcut to look at the coefficient of a, you look at which multiples gives you 'a'. From here you can see that 8a multiply by 4 gives you 32a And 16 multiply by a gives you 16a Since its adding up, 32a+16a=48a

OpenStudy (anonymous):

thank you so it is a^3+12a^2+48a+64

OpenStudy (anonymous):

ok i dont remember using that it has been years so some of this stough i remember how to....

OpenStudy (anonymous):

i feel old brained

OpenStudy (anonymous):

ohhhh k

OpenStudy (anonymous):

now i get it i am so sorry ppl thank you so much for bearing with me i havent been in school fo a long time and i am warming up on old math memories